In 5 minutes, a conveyor belt moves 900 pounds of recyclable aluminum from the delivery truck to a storage area. A smaller belt moves the same quantity of cans the same distance in 7 minutes. If both belts are used, find how long it takes to move the cans to the storage area.
Answer by jorel1380(3719)

1/5+1/7=1/n
7/35+5/35=1/n
12/35=1/n
12n=35
n=35/12 minutes for both belts to move 900 lbs. of aluminum to the storage area..